• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT
    Make New Entry, Make Followup Entry

    User name Sirish Nanda

    Log entry time 20:20:23 on May 1,2010

    Entry number 320402

    keyword=Compton work

    Compton PPLN laser crystal replaced. Swapped in 10W fiber Amp. Fixed optical iolator. Cavity is under turbo vacuum. Final vacuum expected Monday. Locking of the cavity after that.

    NO beam in compton chicane till cavity restoration complete.